### v1

@@ V-E Limitations
The findings of our study are subject to several limitations. First, the evaluation primarily focused on technical development and validation under controlled laboratory conditions. In particular, the validation was limited to the right foot, as the orthosis was consistently worn on the right side. Moreover, the gait analysis was conducted using short treadmill walking sequences, which may not accurately reflect natural gait behavior during prolonged use or in real-world environments. In addition, calibration of the FSR sensors was performed based on manufacturer specifications without subject-specific mechanical calibration, potentially affecting the accuracy of absolute force estimation. Potential variation in sensor placement and the lack of automated recalibration routines may further impact signal reproducibility and long-term stability in real-world scenarios. Furthermore, comparability with treadmill-derived force values was limited, as only aggregated gait data were available from the treadmill, whereas the insole provided full raw data. Finally, the small sample size limits the statistical power of the study and the generalizability of the observed effects.
V-F Future Work
The modular system architecture, which integrates multimodal sensing and feedback, demonstrates how next-generation orthotic devices can move beyond passive monitoring toward interactive and adaptive support. To address current limitations, future work should aim to evaluate the system under real-world conditions and in clinical contexts. A case study with patients affected by gait impairments could provide valuable insights into the system’s robustness, usability, and therapeutic potential.
@@ V-F Future Work
At the application level, the smartphone interface may be extended with features such as chatbots or gamified elements to support user-centered interpretation of health data and increase user engagement. Furthermore, the effectiveness of augmented feedback could be examined in user-centered training scenarios to evaluate its potential impact on motor learning and rehabilitation outcomes. As a next step, we plan to test the orthosis for foot augmentation in virtual reality training environments, with a particular focus on investigating participant feedback interactions in immersive scenarios through visual feedback.
VI Conclusion
@@
Acknowledgment
We thank Diana Völz and Valentin Schwind for their support of our research.

### v2

@@ 5.4 Limitations
The findings of our study are subject to several limitations. First, the evaluation primarily focused on technical development and validation under controlled laboratory conditions. In particular, the validation was limited to the right foot, as the orthosis was consistently worn on the right side. Moreover, the gait analysis was conducted using short treadmill walking sequences, which may not accurately reflect natural gait behavior during prolonged use or in real-world environments. In addition, calibration of the FSR sensors was performed based on manufacturer specifications without subject-specific mechanical calibration, potentially affecting the accuracy of absolute force estimation. Potential variation in sensor placement and the lack of automated recalibration routines may further impact signal reproducibility and long-term stability in real-world scenarios. Furthermore, comparability with treadmill-derived force values was limited, as only aggregated gait data were available from the treadmill, whereas the insole provided full raw data. Similarly, the developed algorithm for gait event detection was not quantitatively compared with existing data-driven gait phase recognition methods, nor validated on publicly available datasets. Finally, the small sample size limits the statistical power of the study and the generalizability of the observed effects, and no control group was included to provide a baseline comparison between participants with and without foot conditions. Nevertheless, the present work was designed as a proof-of-concept investigation with the primary focus on validating the technical feasibility and feedback functionality of the developed sensor system within the intended target group. Consequently, the study focused on end users with foot conditions to ensure practical relevance. Accordingly, the findings represent an initial technical evaluation conducted under controlled laboratory conditions and serve as a foundation for subsequent large-scale studies aimed at generalizing the results.
5.5 Future Work
The modular system architecture, which integrates multimodal sensing and feedback, demonstrates how next-generation orthotic devices can move beyond passive monitoring toward interactive and adaptive support. To address current limitations, future work should aim to evaluate the system under real-world conditions and in clinical contexts. In this context, future testing should include varied environmental conditions, such as uneven outdoor surfaces or stair walking, to evaluate the system’s robustness under realistic gait scenarios. Such experiments will be essential for verifying performance outside controlled laboratory settings. Another direction for future work is the extended testing of the smart insole with regard to material specifications and biocompatibility to ensure long-term durability and safe use in rehabilitation contexts. Specifically, future research should include a larger and more diverse participant sample to allow for more comprehensive statistical analyses and improve the generalizability of the findings. Expanding the sample size will also enable subgroup comparisons and allow the assessment of long-term performance and usability across different user populations. Future studies should further include a control group of participants without foot conditions to establish a baseline for gait parameter comparison and system optimization. Including such a reference group with healthy participants will allow for more comprehensive interpretation of gait-related outcomes and support system optimization. Moreover, a case study involving patients with gait impairments could provide valuable insights into the system’s robustness, usability, and therapeutic potential.
From a methodological perspective, future evaluations should also move beyond subjective usability scales and include behavioral and performance-based assessments, such as latency, energy consumption, correction rate, or changes in gait symmetry when interacting with the feedback system. Additionally, future work could benchmark the algorithm against state-of-the-art gait phase detection approaches and validate it on open-source datasets to enable objective performance comparison and generalization testing.
@@ 5.5 Future Work
At the application level, the smartphone interface may be extended with features such as chatbots or gamified elements to support user-centered interpretation of health data and increase user engagement. Furthermore, the effectiveness of augmented feedback could be examined in user-centered training scenarios to evaluate its potential impact on motor learning and rehabilitation outcomes. As a next step, we plan to test the orthosis for foot augmentation in virtual reality training environments, with a particular focus on investigating participant feedback interactions in immersive scenarios through visual feedback. Additionally, a follow-up study is currently being planned to improve the smartphone application onboarding process using mixed reality, with the goal of enhancing users’ initial interaction, comprehension, and engagement.
6 Conclusions
@@
[…]
Conceptualization, S.R.; methodology, S.R.; software, S.R., A.K., A.C., N.S., F.R., M.Z. and Y.S.; validation, S.R.; formal analysis, S.R.; investigation, S.R.; resources, S.R.; data curation, S.R.; writing—original draft preparation, S.R.; writing—review and editing, S.R. and D.S.-M.; visualization, S.R.; supervision, D.S.-M.; project administration, S.R.; funding acquisition, S.R. All authors have read and agreed to the published version of the manuscript.
[…]
This research was funded by the Hessian Ministry of Science and Art-HMWK, Germany (FL1, Mittelbau), and supported by the RISE program of the German Academic Exchange Service (DAAD).
[…]
The study was conducted in accordance with the Declaration of Helsinki, and approved by the Institutional Ethics Committee of the German Society for Nursing Science (No. 23-027, 12 February 2024).
[…]
Informed consent was obtained from all subjects involved in the study.
[…]
Data are unavailable due to privacy restrictions.
Acknowledgements.
[…]
Appendix A Software Architecture and Data Processing—Smart Orthosis
A.1 Raw Data Acquisition
[…]
A.2 Pre-Processing
[…]
$V_{\text{out}}=\frac{R_{M}\cdot V^{+}}{R_{M}+R_{\text{FSR}}}$ (1)
Following the manufacturer’s specification Interlink Electronics (2013), the raw sensor output was used to estimate relative plantar forces without applying sensor-specific calibration. To reduce high-frequency noise, a moving average filter was applied to the estimated force signals (Equation (2)).
$\text{filteredForce}[i]=\frac{1}{N}\sum_{k=0}^{N-1}\text{fsrForce}[i-k]$ (2)
The BMI270 includes internal configuration options for filtering and calibration, which were utilized to enhance signal quality. The sensor was configured with a measurement range of $\pm 4$ g for the accelerometer and $\pm 2000^{\circ}/$s for the gyroscope to capture typical gait-related movements as well as rapid rotational changes. Offset correction was automatically performed during initialization by fixing the IMU in a stationary position to determine the static bias for each axis. These offset values were removed from all subsequent accelerometer and gyroscope samples to ensure base-corrected measurements (Equations (3) and (4)).
$\text{accelCorrected}[i]=\text{accelRaw}[i]-\text{accelOffset}$ (3)
$\text{gyroCorrected}[i]=\text{gyroRaw}[i]-\text{gyroOffset}$ (4)
Additionally, separate internal low-pass filters were applied to the accelerometer and gyroscope axes to remove high-frequency noise (Equation (5)).
$\text{filteredValue}[i]=\alpha\cdot\text{filteredValue}[i-1]+(1-\alpha)\cdot\text{rawValue}[i]$ (5)
A filter coefficient of $\alpha=0.9$ was chosen to effectively reduce high-frequency noise. Pre-processing of the IMU data through offset correction and low-pass filtering ensures that the acceleration and rotation data are precise and reliable.
A.3 Signal Processing
In the signal processing stage, pre-processed sensor data were further refined to reduce residual noise and enable robust detection of gait events. To improve signal quality, Kalman filters Kalman (1960) were applied to both FSR and IMU signals. The Kalman filter was selected due to its proven suitability for real-time gait analysis in wearable systems, including IMU-based orientation estimation, walking distance calculation, and general gait assessment Prasanth et al. (2021); Shankar et al. (2010); Bennett et al. (2013). For the FSR signals, individual Kalman filters were assigned to each of the sensor channels, with parameters set to $Q=1.0$, $R=1.0$, and $P=0.01$. The filter parameters were adjusted to maintain responsiveness while reducing high-frequency fluctuations. For the IMU, each axis of the accelerometer and gyroscope data was filtered individually, with configuration parameters adjusted to the sensor’s noise characteristics. Filter parameters were empirically set to $Q=0.1$, $R=0.1$, and $P=0.3$ for accelerometer data, and to $Q=1.0$, $R=1.0$, and $P=0.01$ for gyroscope data to balance signal smoothness and responsiveness.
The filtered FSR signals served as the basis for gait event detection. A custom rule-based algorithm was implemented to identify key gait phases using threshold values defined as relative percentages of the user’s body weight. The algorithm evaluated pressure variations within specific sensor zones corresponding to the forefoot and hindfoot regions to detect events such as heel strike, toe strike, heel off, and toe off. This individualized approach enabled the detection of characteristic loading patterns and supported the identification of critical gait events and potential overload conditions. Thresholds were empirically determined through pre-testing across multiple gait cycles and optimized to minimize detection errors. A finite state machine was implemented to manage gait phase transitions based on these events. This approach ensured a consistent event sequence and enabled accurate temporal segmentation. A baseline correction algorithm was used to counteract the hysteresis effects of the FSR sensors and to ensure zero-base alignment after the load is removed. This adjustment was essential for maintaining continuous and reliable gait event detection across steps. The detected gait events, including their timestamps, were subsequently used to compute spatiotemporal gait parameters such as step duration, stance/swing phase duration, and cadence.
A.4 Sensor Fusion
Filtered FSR and IMU data were fused to extract both temporal and spatial characteristics of the gait cycle. While FSR data provide reliable information on foot-ground contact timing, the IMU captures the foot’s three-dimensional orientation and motion dynamics. A Mahony filter Mahony et al. (2008) was employed to estimate foot orientation in terms of roll, pitch, and yaw. The Mahony filter was implemented via the fusion.MahonyUpdate() function from the SensorFusion library (2023) (aster94). This approach combines gyroscope data for short-term tracking with accelerometer data for long-term correction, effectively compensating for drift and high-frequency noise. To improve the accuracy of position estimation, zero-velocity updates Skog et al. (2010) were applied during the stance phase, identified from FSR signals. These updates reset the velocity to zero between the heel contact and the heel off, thereby limiting integration drift and reducing cumulative error in position calculation.
A.5 Extraction of Gait Parameters
Based on the fused sensor data, key gait parameters were extracted. The segmentation of the gait cycle follows the definition of Perry Perry et al. (2010), which distinguishes two main phases: the stance phase (60% of the gait cycle) and the swing phase (40%), as well as eight sub-phases (including foot strike, foot flat, heel off, and foot off). These gait parameters can be classified into spatial, temporal, spatio-temporal, and kinetic categories Lovecchio et al. (2021). Stride length was calculated by integrating acceleration data, corrected for foot orientation and stabilized using zero-velocity updates during stance phases. Gait speed was subsequently derived as the ratio of stride (cycle) length to gait cycle duration. As the measurement setup included only a single foot-mounted IMU, step length and step width could not be directly calculated. However, when the instrumented insoles are worn on both feet, these parameters can be extracted. In addition to standard parameters, a stability index was introduced to quantify gait stability. This metric integrates two normalized components:
[…]